OpenStudy (anonymous):

Five of the six exterior angle measures of a nonregular hexagon measure 55°, 60°, 69°, 57°, and 57°. What is the measure of the interior angle adjacent to the sixth exterior angle?

OpenStudy (anonymous):

*exterior angle of every polygon measures 360 so 55+60+69+57+57+x=360 298+x=360 x=360-298 x=62 *the interior angle of a polygon is find using the formula "[n-2]180"n is the side of the polygon [n-2]180 [6-2]180 4 x180 interior angle=720